### v1

@@ 3.3.2 Curve Fitting
Both curve fitting and MCMC fitting in Section 3.3.3 are done in a narrower region called the “focus region”, which is dynamically determined according to the central wavelength of the DIB pre-detected in Section 3.3.1. The flux values beyond the focus region are set to the continuum level to prevent the other potential features, e.g. certain weaker DIBs (see the bottom panel of Figure 3), from disturbing the fitting. The wavelength interval of the focus region is designed based on 2-$\sigma$ principle, where $\sigma$ refers to the Gaussian width. For instance, if a DIB has a $\sigma$ of 2 Å, its entire profile would span approximately $2\times 3\sigma$, i.e., 12 Å. Therefore, a wavelength interval of at least $82\times 2\sigma=$ 8 Å is necessary for a reasonable fitting. Specifically, for the DIBs $\lambda 5797$ and $\lambda 6614$, the width of the focus region is set to 6 Å on both sides of the pre-detected central wavelength. For the DIB $\lambda 5780$, the width is set to 10 Å because the typical FWHM of the DIB $\lambda 5780$ is larger than those of the DIBs $\lambda 5797$ and $\lambda 6614$ (Hobbs et al. 2008; Fan et al. 2019). On top of the above cases, if the DIBs $\lambda 5780$ and $\lambda 5797$ are both pre-detected, the blue end of the focus region is set to 6 Å on the left side of the pre-detected central wavelength of the DIB $\lambda 5780$, and the red end is set to 6 Å on the right side of the pre-detected central wavelength of the DIB $\lambda 5797$. Similar to the wavelength interval of the pre-detection in Section 3.3.1, such a focus region is wide enough to handle the broadened DIBs, which may be caused by multiple DIB components or the instrumental broadening effect.
We employ the method of least-square curve fitting (the function curve_fit in the Python module scipy.optimize) to obtain the central wavelength, depth and FWHM of the single Gaussian profile of the DIB as Equation 3.
@@
where $x$ means the wavelength, and $\rm D$, $\rm\mu$, $\rm\sigma$ denote the depth, central wavelength and Gaussian width, respectively. In the case where the DIBs $\lambda 5780$ and $\lambda 5797$ are both detected, a double Gaussian profile is adopted to simultaneously fit the two DIBs, which can overcome the blend of the two close DIBs when the DIBs are greatly broadened. The initial values of the depth and central wavelength are taken from the pre-detection results in Section 3.3.1, and the initial values of the Gaussian width are set to 3.85, 2 and 2 Å for the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$, respectively. Such initial values are larger than the typical values of the DIBs owing to the instrumental broadening effect in the low resolution (Yuan & Liu 2012; Lan et al. 2015). Actually, there is little influence on the final fitting results while using the canonical Gaussian width, e.g., the FWHM from Fan et al. (2019), as the initial value.
@@ 3.3.3 MCMC Fitting
The prior in the Bayesian inference can rule out some cases that obviously violate the physics, e.g., the central wavelength of DIB is far away from the well-known value. And the MCMC method is able to efficiently sample the multi-parameter distribution from the posterior constructed by the product of the likelihood and prior based on the Bayesian theory (Equation 4).
@@
where $\rm\theta$ represents the parameters of the DIB profile, i.e., $\rm\theta=\left(D,\mu,\sigma\right)$, and $\rm Y$ denotes the observed data. Specifically, the likelihood and prior are expressed as Equations 5 and 6 in logarithmic form, respectively.
@@
where $x$, $y$ and $y_{\text{err}}$ as the observed data $\rm Y$ are the wavelength, flux and flux error of the ISM residual spectrum, respectively, and $f_{\theta}(x)$ is the model of DIB profile which is set out in Equation 3.
@@
where $\rm P(D)$, $\rm P(\mu)$ and $\rm P(\sigma)$ are the priors of the depth, central wavelength and Gaussian width, respectively. The prior of $\rm D$ satisfies the uniform distribution, and the priors of $\rm\mu$ and $\rm\sigma$ are both set to be the Gaussian distributions with the means of the curve fitting results in Section 3.3.2 and the standard deviations of 0.5 Å. Table 2 summarizes the detailed upper and lower limits of the priors of $\rm D$, $\rm\mu$ and $\rm\sigma$ for the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$.
@@
\$1\$\$1\$footnotetext: The upper limit of the prior of $\rm D$ is set according to the maximum depth of all the targets measured by pre-detection in Section 3.3.1.
The Python package emcee is called to carry out the MCMC fitting. We set 100 walkers and 250 steps for each walker, where the first 50 steps are used as the burn-in stage and the last 200 steps are used to sample the posterior distribution. The initial guesses of the walkers are randomly seeded around the curve fitting results in Section 3.3.2 with a standard deviation of 0.01, which can greatly reduce the length of the chain as a result of the fast convergence of the MCMC fitting. We take the 50th, 16th and 84th percentiles of the posterior distribution sampled by the MCMC as our best estimate and statistical uncertainty. The equivalent width (EW) of the DIB is calculated by $\rm EW=\sqrt{2\pi}\times D\times\sigma$. The EW error is estimated by the error propagation from the errors of $\rm D$ and $\rm\sigma$.
@@ 4.1 Quality Control
4.1.1 Define the High Quality Samples
A total of 2,188,240 target spectra are processed by the pipeline described in Section 3, and the statistics of all the valid measurements for the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$ are summarized in Table 3. To analyze the results based on the reliable DIB measurements, we define the high quality (HQ) samples. For each HQ sample, it must meet the following criteria:
@@ 4.1.1 Define the High Quality Samples
    The depth fitted by the MCMC $\rm D_{mc}>$ 3 / SNR_r. The threshold of 3 / SNR_r is set to ensure the DIB signal is significantly above the noise level.
@@
    The coefficient of variation (CV) of the depths $\rm\{D_{pr},\ D_{cf},\ D_{mc}\}<$ 10%, where the CV is defined as the standard deviation divided by the mean value, and the subscripts $\rm pr$, $\rm cf$ and $\rm mc$ represent the pre-detection, curve fitting and MCMC fitting in Section 3, respectively.
@@
    The Gaussian width $\rm\sigma_{mc}>$ 1 Å and $\rm\sigma_{mc}<$ 3 Å, where the lower limit of 1 Å is set due to the instrumental broadening effect (see Appendix A), and the upper limit of 3 Å is set based on the typical FWHM of DIB (Vogrinčič et al. 2023), which will be elaborated in Section 4.2.1.
@@
    $\rm EW_{DIB}/EW_{cont}>$ 33% for the DIB $\lambda 5780$ and $>$ 25% for the DIBs $\lambda 5797$ and $\lambda 6614$, where $\rm EW_{DIB}$ is the equivalent width of the DIB, and $\rm EW_{cont}$ is the integral equivalent width of the continuum outside the DIB region, i.e., $\rm[\mu_{mc}-3\sigma_{mc},\ \mu_{mc}+3\sigma_{mc}]$. The constraints can ensure the DIBs rather than the stellar residuals or noises significantly contribute to the EWs.
Besides the above criteria, we also tighten the constraints on the CV and central wavelength for the M- and K-type ($T_{\rm eff}<5200\ \rm K$) stars to get rid of the fake DIB signals caused by the stronger metal absorptions in cooler stars. Specifically, the CV must be less than 5% and the standard deviation of $\rm\{\mu_{pr},\ \mu_{cf},\ \mu_{mc}\}$ must be less than 0.25 Å. After applying the above criteria, we obtain the HQ samples, and the summary of the HQ samples are also listed in Table 3. There are a total of 7,681 stars whose spectra simultaneously contain the HQ measurements of the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$. Besides the public access to three catalogs of the HQ samples for the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$, here we also release the results of these 7,681 HQ sources  [^7] We emphasize that the “HQ sample” or “HQ measurement” means the individual DIB measurement of each cool stellar spectrum, and the “HQ source” is on behalf of the mean value of the DIB measurements of the multiple-epoch spectra for the same star. See more details in Section 4.2.3 and Section 5. via https://nadc.china-vo.org/res/r101404/.
4.1.2 Define the Marginal Quality Samples
Aside from the HQ samples, the marginal quality (MQ) samples are also available online. Although the MQ samples can not provide DIB measurements as reliable as the HQ samples, these low or null values can offer important boundary constraints for other studies, e.g., 3D mapping. For each MQ DIB measurement, it is required to meet the following two criteria.
@@ 4.1.2 Define the Marginal Quality Samples
    The EW fitted by the MCMC, $\rm EW_{DIB}\leq 0.03$, where the upper limit is set to 0.03 Å according to the typical uncertainty of 0.04 Å for the DIB measurements validated in Section 4.2.2, and the lowest EW of around 0.03 Å measured in the HQ samples.
@@
    $\rm EW_{cont}\leq 0.5$, where $\rm EW_{cont}$ is the integral equivalent width of the continuum outside the DIB region, i.e., $\rm[\mu_{mc}-3\sigma_{mc},\ \mu_{mc}+3\sigma_{mc}]$. The upper limit is set to 0.5 Å to ensure that the stellar residuals are cleaned up to the continuum level.
After applying the above criteria, there are a total of 27,598, 42,735 and 21,635 MQ measurements of the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$, respectively, corresponding to 23,920, 36,825 and 18,145 unique stars. Note that we do not set a specific threshold to further distinguish between marginal and null values, as the uncertainties estimated by MCMC are provided in the catalogs. Therefore, Users can determine the threshold based on their needs. However, we suggest using a ratio of $\rm err(EW_{DIB})/\rm EW_{DIB}$ of 1 as the dividing line: values below 1 can be considered marginal, while values above 1 can be classified as null.
@@ 4.2 Validation
4.2.1 Central Wavelength V.S. Gaussian Width
The $\lambda-\sigma$ 2D-histogram has been a powerful diagnostic tool to visually distinguish the real DIBs from the fake ones in the context of automation measurement of DIB for a large spectroscopic survey (Saydjari et al. 2023; Zhao et al. 2024). The fake DIBs which may be the stellar absorptions are usually far away from the typical central wavelength of the DIBs, and their Gaussian widths are narrower than the typical values of the real DIBs as the FWHM of DIB is generally broader than that of metal absorption line. In the 2D-histogram, the true DIB measurements are located around the rest-frame central wavelength and typical Gaussian width of the DIB, but with certain scatter, whereas the wrong measurements or stellar features exhibit small Gaussian widths and are highly concentrated in a region.
@@
Figure 5: Two-dimensional histograms of the measured DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$ as a function of the Gaussian width $\sigma$ and the central wavelength $\lambda$ in the rest frame. (a) The three upper panels present the density of all the samples that may have DIBs, (b) and the other three lower panels show the density of the HQ samples which have the reliable DIB measurements. All of the histograms have the same bin size of 100. The red horizontal dashed lines denote the central wavelengths of the DIBs, which are fitted by the method in Section 5.3. The red vertical dashed lines indicate the typical Gaussian widths of the DIBs considering the instrumental broadening of LAMOST LRS (see Appendix A), which are taken from Vogrinčič et al. (2023).
4.2.2 Integral EW V.S. Fitted EW
To quantify the reliability of the DIB measurements, we compute the integral EW ($\rm EW_{int}$) of the three DIBs in the HQ samples within the same wavelength range, i.e., $\rm\mu_{mc}\pm 3\sigma_{mc}$ as the fitted EW ($\rm EW_{fit}$). As shown in Figure 6, the $\rm EW_{int}$ is highly consistent with the $\rm EW_{fit}$, and the difference between them ($\rm\Delta EW$) is mainly concentrated in a tiny region that is less than 0.04 Å for over 99% of the samples among three DIBs. For all three DIBs, however, there are certain samples whose $\rm EW_{int}$ are far away from their $\rm EW_{fit}$, which stem from the fact that the continuum level is not well determined (Farhang et al. 2015b; Farhang et al. 2019). For the DIB $\lambda 5780$, the blend with a very broad DIB $\lambda 5778$ is a confounding factor in the determination of the continuum level (Herbig 1975; Krelowski & Walker 1987; Herbig 1993). Similar with the DIB $\lambda 5780$, the blue wing of the DIB $\lambda 5797$ is also overlapped with a quite broad DIB $\lambda 5795$ (Krelowski & Walker 1987; Friedman et al. 2011). Nevertheless, Vos et al. (2011) has reported that the broader DIBs $\lambda 5778$ and $\lambda 5795$ are largely eliminated from the ISM residual spectra in the local normalization, and they are too weak to significantly contaminate the EW measurements of the DIBs $\lambda 5780$ and $\lambda 5797$, even for high extinction. Such broader DIB blend does not exist for the DIB $\lambda 6614$ (Herbig & Soderblom 1982; Galazutdinov et al. 2008), which gives rise to better agreement (see Figure 6(c)) between the $\rm EW_{int}$ and the $\rm EW_{fit}$ than the other two DIBs.

### v2

@@ 3.3.2 Curve fitting
[…]
We employ the method of least-square curve fitting (the function curve_fit in the Python module scipy.optimize) to obtain the central wavelength, depth and FWHM of the single Gaussian profile of the DIB as Eq. (3).
@@
where $x$ means the wavelength, and $\rm D$, $\rm\mu$, $\rm\sigma$ denote the depth, central wavelength, and Gaussian width, respectively. In the case where the DIBs $\lambda 5780$ and $\lambda 5797$ are both detected, a double Gaussian profile is adopted to simultaneously fit the two DIBs, which can overcome the blend of the two close DIBs when the DIBs are greatly broadened. The initial values of the depth and central wavelength are taken from the pre-detection results in Sect. 3.3.1, and the initial values of the Gaussian width are set to 3.85, 2, and 2 Å for the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$, respectively. Such initial values are larger than the typical values of the DIBs owing to the instrumental broadening effect in the low resolution (Yuan & Liu 2012; Lan et al. 2015). Actually, there is little influence on the final fitting results while using the canonical Gaussian width, for instance, the FWHM from Fan et al. (2019), as the initial value.
@@ 3.3.3 MCMC fitting
The prior in the Bayesian inference can rule out some cases that obviously violate the physics, for example, the central wavelength of DIB is far away from the well-known value. And the MCMC method is able to efficiently sample the multi-parameter distribution from the posterior constructed by the product of the likelihood and prior based on the Bayesian theory (Eq. (4)).
@@
where $\rm\theta$ represents the parameters of the DIB profile, that is, $\rm\theta=\left(D,\mu,\sigma\right)$, and $\rm Y$ denotes the observed data. Specifically, the likelihood and prior are expressed as Eqs. (5) and (6) in logarithmic form, respectively.
@@
where $x$, $y$, and $y_{\text{err}}$ as the observed data $\rm Y$ are the wavelength, flux, and flux error of the ISM residual spectrum, respectively, and $f_{\theta}(x)$ is the model of DIB profile which is set out in Eq. (3).
@@
where $\rm P(D)$, $\rm P(\mu)$, and $\rm P(\sigma)$ are the priors of the depth, central wavelength, and Gaussian width, respectively. The prior of $\rm D$ satisfies the uniform distribution, and the priors of $\rm\mu$ and $\rm\sigma$ are both set to be the Gaussian distributions with the means of the curve fitting results in Sect. 3.3.2 and the standard deviations of 0.5 Å. Table 2 summarizes the detailed upper and lower limits of the priors of $\rm D$, $\rm\mu$, and $\rm\sigma$ for the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$.
@@
\$1\$\$1\$footnotetext: The upper limit of the prior of $\rm D$ is set according to the maximum depth of all the targets measured by pre-detection in Sect. 3.3.1.
The Python package emcee is called to carry out the MCMC fitting. We set 100 walkers and 250 steps for each walker, where the first 50 steps are used as the burn-in stage and the last 200 steps are used to sample the posterior distribution. The initial guesses of the walkers are randomly seeded around the curve fitting results in Sect. 3.3.2 with a standard deviation of 0.01, which can greatly reduce the length of the chain as a result of the fast convergence of the MCMC fitting. We take the 50th percentile and half of the difference between the 16th and the 84th percentiles of the posterior distribution sampled by the MCMC as our best estimate and statistical uncertainty. The equivalent width (EW) of the DIB is calculated by $\rm EW=\sqrt{2\pi}\times D\times\sigma$. The EW error is estimated by the error propagation from the errors of $\rm D$ and $\rm\sigma$.
@@ 4.1 Quality control
4.1.1 Defining the high-quality samples
A total of 2,188,240 target spectra are processed by the pipeline described in Sect. 3, and the statistics of all the valid measurements for the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$ are summarized in Table 3. To analyze the results based on the reliable DIB measurements, we define the high-quality (HQ) samples. For each HQ sample, it must meet the following criteria:
@@ 4.1.1 Defining the high-quality samples
    The depth fitted by the MCMC $\rm D_{mc}>\frac{3}{S/N}$. The threshold of $\rm\frac{3}{S/N}$ is set to ensure the DIB signal is significantly above the noise level.
@@
    The coefficient of variation (CV) of the depths $\rm\{D_{pr},\ D_{cf},\ D_{mc}\}<$ 10%, where the CV is defined as the standard deviation divided by the mean value, and the subscripts $\rm pr$, $\rm cf$ and $\rm mc$ represent the pre-detection, curve fitting and MCMC fitting in Sect. 3, respectively.
@@
    The Gaussian width $\rm\sigma_{mc}>$ 1 Å and $\rm\sigma_{mc}<$ 3 Å, where the lower limit of 1 Å is set due to the instrumental broadening effect (see Appendix A), and the upper limit of 3 Å is set based on the typical FWHM of DIB (Vogrinčič et al. 2023), which will be elaborated in Sect. 4.2.1.
@@
    $\rm EW_{DIB}/EW_{cont}>$ 33% for the DIB $\lambda 5780$ and $>$ 25% for the DIBs $\lambda 5797$ and $\lambda 6614$, where $\rm EW_{DIB}$ is the equivalent width of the DIB, and $\rm EW_{cont}$ is the integral equivalent width of the continuum outside the DIB region, namely $\rm[\mu_{mc}-3\sigma_{mc},\ \mu_{mc}+3\sigma_{mc}]$. The constraints can ensure the DIBs rather than the stellar residuals or noises significantly contribute to the EWs.
Besides the above criteria, we also tighten the constraints on the CV and central wavelength for the M- and K-type ($T_{\rm eff}<5200\ \rm K$) stars to get rid of the fake DIB signals caused by the stronger metal absorptions in cooler stars. Specifically, the CV must be less than 5% and the standard deviation of $\rm\{\mu_{pr},\ \mu_{cf},\ \mu_{mc}\}$ must be less than 0.25 Å. After applying the above criteria, we obtain the HQ samples, and the summary of the HQ samples are also listed in Table 3. There are a total of 7,681 spectra simultaneously contain the HQ measurements of the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$. Besides the public access to three catalogs of the HQ samples for the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$, here we also release the results of these 7,681 HQ measurements  [^7] We emphasize that the “HQ sample” or “HQ measurement” refers to the individual DIB measurement of each cool stellar spectrum, and the “HQ source” refers to the mean value of the DIB measurements of the multiple-epoch spectra for the same star. See more details in Sects. 4.2.3 and 5. via https://nadc.china-vo.org/res/r101404/.
4.1.2 Defining the marginal-quality samples
Aside from the HQ samples, the marginal-quality (MQ) samples are also available online. Although the MQ samples cannot provide DIB measurements as reliable as the HQ samples, these low or null values can offer important boundary constraints for other studies, for example, 3D mapping. For each MQ DIB measurement, it is required to meet the following two criteria.
@@ 4.1.2 Defining the marginal-quality samples
    The EW fitted by the MCMC, $\rm EW_{DIB}\leq 0.03$, where the upper limit is set to 0.03 Å according to the typical uncertainty of 0.04 Å for the DIB measurements validated in Sect. 4.2.2, and the lowest EW of around 0.03 Å measured in the HQ samples.
@@
    $\rm EW_{cont}\leq 0.5$, where $\rm EW_{cont}$ is the integral equivalent width of the continuum outside the DIB region, namely $\rm[\mu_{mc}-3\sigma_{mc},\ \mu_{mc}+3\sigma_{mc}]$. The upper limit is set to 0.5 Å to ensure that the stellar residuals are cleaned up to the continuum level.
After applying the above criteria, there are a total of 27,598, 42,735, and 21,635 MQ measurements of the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$, respectively, corresponding to 23,920, 36,825, and 18,145 unique stars. We note that we do not set a specific threshold to further distinguish between marginal and null values, as the uncertainties estimated by MCMC are provided in the catalogs. Therefore, users can determine the threshold based on their needs. However, we suggest using a ratio of $\rm err(EW_{DIB})/\rm EW_{DIB}$ of 1 as the dividing line: values below 1 can be considered marginal, while values above 1 can be classified as null.
@@ 4.2 Validation
4.2.1 Central wavelength versus Gaussian width
The $\lambda-\sigma$ 2D histogram has been a powerful diagnostic tool for visually distinguishing real DIBs from false positives in the context of automatic measurement of DIBs from a large spectroscopic survey (Saydjari et al. 2023; Zhao et al. 2024). The false-positive DIBs, which may be stellar absorption, are usually far away from the typical central wavelength of the DIBs, and their Gaussian widths are narrower than the typical values of the real DIBs as the FWHM of a DIB is generally broader than that of metal absorption line. In the 2D histogram, the true DIB measurements are located around the rest-frame central wavelength and typical Gaussian width of the DIB, but with a certain scatter, whereas the stellar features exhibit smaller Gaussian widths and are highly concentrated in a region.
@@ 4.2.1 Central wavelength versus Gaussian width
Figure 5: Two-dimensional histograms of the measured DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$ as a function of the Gaussian width $\sigma$ and the central wavelength $\lambda$ in the rest frame. (a) The three upper panels present the density of all the samples that may have DIBs, (b) and the other three lower panels show the density of the HQ samples for which we have reliable DIB measurements. All of the histograms have the same bin size of 100. The red horizontal dashed lines denote the central wavelengths of the DIBs, which are fitted using the method described in Sect. 5.3. The red vertical dashed lines indicate the typical Gaussian widths of the DIBs considering the instrumental broadening of LAMOST LRS (see Appendix A), which are taken from Vogrinčič et al. (2023).
4.2.2 Integral EW versus fitted EW
To quantify the reliability of the DIB measurements, we compute the integral EW ($\rm EW_{int}$) of the three DIBs in the HQ samples within the same wavelength range, that is, $\rm\mu_{mc}\pm 3\sigma_{mc}$ as the fitted EW ($\rm EW_{fit}$). As shown in Fig. 6, the $\rm EW_{int}$ is highly consistent with the $\rm EW_{fit}$, and the difference between them ($\rm\Delta EW$) is mainly concentrated in a tiny region that is less than 0.04 Å for over 99% of the samples among three DIBs. For all three DIBs, however, there are certain samples whose $\rm EW_{int}$ are far away from their $\rm EW_{fit}$, which stem from the fact that the continuum level is not well determined (Farhang et al. 2015b; Farhang et al. 2019). For the DIB $\lambda 5780$, the blend with a very broad DIB $\lambda 5778$ is a confounding factor in the determination of the continuum level (Herbig 1975; Krelowski & Walker 1987; Herbig 1993). Similar with the DIB $\lambda 5780$, the blue wing of the DIB $\lambda 5797$ is also overlapped with a quite broad DIB $\lambda 5795$ (Krelowski & Walker 1987; Friedman et al. 2011). Nevertheless, Vos et al. (2011) has reported that the broader DIBs $\lambda 5778$ and $\lambda 5795$ are largely eliminated from the ISM residual spectra in the local normalization, and they are too weak to significantly contaminate the EW measurements of the DIBs $\lambda 5780$ and $\lambda 5797$, even for high extinction. Such broader DIB blend does not exist for the DIB $\lambda 6614$ (Herbig & Soderblom 1982; Galazutdinov et al. 2008), which gives rise to better agreement (see Fig. 6(c)) between the $\rm EW_{int}$ and the $\rm EW_{fit}$ than the other two DIBs.